OpenStudy (anonymous):

Rachel bought 6 1/4 pounds of dried fruit. If 1/3 of the fruit was dried apricots, how many pounds of dried apricots did she buy? 6 1/12 2 1/12 2 1/4 2 1/6

OpenStudy (anonymous):

Please explain

OpenStudy (anonymous):

Simply find a third of 6 1/4. You do not need to ask for explanation. I would be glad to provide it, even if you did not ask for it ;) Anyway, To make the first composite fraction as a single fraction, the step is to Multiply the denominator by the whole number and add with the numerator. the result of this will be the numerator of the new fraction and the denominator will remain the same, meaning:\[6\frac{ 1 }{ 4 }= \frac{ (4*6)+1 }{ 4 }= \frac{ 25 }{ 4 }\]

OpenStudy (anonymous):

Next, you find a third of those pounds to be those of dried apricots. To do this, you multiply the first fraction by the inverse of the second, giving us\[\frac{ 25 }{ 4}*\frac{ 3 }{ 1 }=\frac{ 75 }{ 4 }\]

OpenStudy (anonymous):

As I explained in how to change from composite fraction to simple fraction, you can do that again, by seeing which one of these fraction give 75/4 and it is

OpenStudy (anonymous):

ok thank you :3

OpenStudy (anonymous):

Wait. in my second explanation, the fraction should have been \[\frac{ 25 }{ 4 }*\frac{ 1 }{ 3 }=\frac{ 25 }{ 12 }\]

OpenStudy (anonymous):

And the answer becomes B.

OpenStudy (anonymous):

0.0 k..

OpenStudy (anonymous):

thx it says 100% on answer thinggy

OpenStudy (anonymous):

Medal please?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

der have a awesome day :D

OpenStudy (anonymous):

if i could i hire u as my tutor >.<

OpenStudy (anonymous):

You have a nice and twice as bright of an usual day!

OpenStudy (anonymous):

That is no problem. You can always tag my name and I will be able to help whenever possible or we could email.

OpenStudy (anonymous):

ok :p